Given a right triangle, find the measures of all the angles, in degrees, if one angle is a right angle and the measure of the second angle is six less than seven times the measure of the third angle.
Found 2 solutions by VFBundy, ikleyn:
Answer by VFBundy(438)

First angle = 90
Second angle = x
Third angle = 90 - x

x = 7(90 - x) - 6

x + 6 = 7(90 - x)

x + 6 = 630 - 7x

8x = 624

x = 78

First angle = 90
Second angle = x = 78
Third angle = 90 - x = 90 - 78 = 12

Answer by ikleyn(52786)

.

When (and since) you called a triangle as a right triangle,

you should not explain in the next phrase that one of its angles is the right angle.


You just defined it once ------ it is just enough.